OpenStudy (anonymous):

Simplify the expression 5^-3.

OpenStudy (lilg132):

1/125

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

= 1 / 5^3 = 1 / 125

OpenStudy (lilg132):

np